# BALF: Budgeted Activation-Aware Low-Rank Factorization for Fine-Tuning-Free Model Compression

### v1

@@ 4.4 Putting It All Together: BALF
6:  ${\bm{M}},{\bm{M}}^{+}]\leftarrow\textsc{BatchedUWM}(\overline{I}_{l}({\bm{\mathsfit{X}}}^{l-1})^{T}\overline{I}_{l}({\bm{\mathsfit{X}}}^{l-1}));\ {\bm{\mathsfit{W}}}\leftarrow\mathcal{M}_{l}.\operatorname{weights}$
@@
13:  ${\bm{U}},{\bm{\Sigma}},{\bm{V}}^{T},{\bm{M}},{\bm{M}}^{+}]\leftarrow F[l]$; ${\bm{W}}_{0}\leftarrow{\bm{M}}{\bm{U}}{\bm{\Sigma}}$; ${\bm{W}}_{1}\leftarrow{\bm{V}}^{T}$
@@
A detailed statement of the result, its proof, and special cases appear in Appendix E. The bound’s quality depends on model depth and properties of the weights, and it is typically loose in practical deep networks. Even so, it remains conceptually informative.
@@ 5 Experiments
First, we perform an extensive sweep showcasing the complexity-accuracy curves for FLOP and parameter counts. To emphasize what parts of BALF are important, each run is performed with different methods—SVD with the energy criterion, activation-aware SVD with the energy criterion, and BALF (activation-aware decomposition with our rank allocator); see Section 4.3 for an overview.
@@
For ResNet-50, we compare BALF with several pruning methods: DFPC (Narshana et al., 2023), IterTVSPrune (ITVSP) (Murti et al., 2023) and IFM (Chen et al., 2024). For ResNet-20, we compare against the pruning approach FVRCP (He et al., 2021) and the factorization method ALDS (Liebenwein et al., 2021).
@@
Compressing compact models such as MobileNet-V2 is known to be challenging. For this model, we compare against the factorization methods SVD-NAS (Yu & Bouganis, 2022), ALDS (Liebenwein et al., 2021), and LR-S2 (Idelbayev & Carreira-Perpiñán, 2020) (using the no-fine-tuning results for ALDS and LR-S2 reported by Yu & Bouganis (2022)). We also compare with the pruning-restoration method proposed by Lee et al. (2025), using their L2 variant (denoted L2+REST). Note that SVD-NAS requires an expensive search procedure (on the order of hours), and LR-S2 allocates ranks through a training run, making them substantially more expensive than BALF. Table 1 contains the different results. In general, BALF surpasses or is competitive with all the compared methods.
@@
On an RTX 2070 laptop, compressing ImageNet models completes in minutes (under four minutes for all models except ResNeXt-101, taking less than seven). The most expensive steps can be cached, reducing subsequent compression runs to seconds. We provide a comprehensive benchmark and implementation details in Appendix H.
@@
BALF-F-0.5 -50.04 -43.86 -13.84
BALF-P-0.5 -28.18 -50.08 -7.78
BALF-F-0.7 -30.00 -21.69 -3.48
BALF-P-0.7 -13.55 -29.94 -1.96
BALF-F-0.8 -20.00 -9.77 -1.65
BALF-P-0.8 -9.03 -21.92 -1.12
IFM -5.65 -1.46
IFM -20.42 -10.45
ITVSP -4.76 -3.02
ITVSP -9.98 -10.21
DFPC -5.66 -5.78
DFPC -10.92 -13.88
@@
BALF-F-0.5 -50.02 -49.48 -24.89
BALF-P-0.5 -24.75 -49.74 -13.21
BALF-F-0.7 -30.74 -29.10 -6.26
BALF-P-0.7 -12.77 -29.88 -2.50
BALF-F-0.8 -20.03 -17.90 -2.16
BALF-P-0.8 -9.27 -21.70 -1.20
ALDS -23.71 -1.50
ALDS -30.05 -3.37
ALDS -50.00 -25.22
FVRCP -19.84 -2.5
FVRCP -29.51 -5.83
FVRCP -49.35 -24.86
@@
BALF-F-0.6 -40.00 -39.52 -3.88
BALF-P-0.6 -40.36 -39.87 -4.11
BALF-F-0.7 -30.00 -29.64 -1.08
BALF-P-0.7 -30.27 -29.91 -1.09
BALF-F-0.8 -20.03 -19.79 -0.14
BALF-P-0.8 -20.18 -19.94 -0.16
GTP-15.3 -13.07 0.00 +0.00
GTP-8.8 -50.00 0.00 -3.50
PRACTISE -16.6 -2.5
DC-ViT -16.6 -0.54
@@
BALF-P-0.7 -2.06 -29.70 -27.65
BALF-P-0.75 -1.11 -24.79 -6.13
BALF-P-0.8 -0.23 -19.80 -0.59
BALF-P-0.9 -0.13 -10.76 -0.25
BALF-F-0.97 -2.51 -1.55 -4.95
SVD-NAS -15.09 -9.00 -12.54
ALDS -2.62 -37.61 -16.95
LR-S2 -3.81 -6.24 -17.46
L2+REST -5.00 -4.42
L2+REST -10.00 -18.47
@@
On CIFAR-10-C (Hendrycks & Dietterich, 2019), BALF remained robust to distribution shift: when whitening matrices were estimated on the standard CIFAR-10 training set and evaluation was performed on CIFAR-10-C, the additional accuracy drop was typically less than 2.5 pp. Moreover, we found that, often, a smaller number of calibration samples (about 256 and 1024 for CIFAR-10 and ImageNet, respectively) yields only marginally worse results than those reported in this section. For instance, see Figure 3 for a sweep on ResNet-50. The remaining results can be found in Appendix G.
@@
Even when surpassing (or being competitive with) other methods, compressing compact models (like MobileNet-V2) remains difficult. End-to-end speedups are notable, but we believe implementing specialized operators would yield better gains (see Section G.3 for a discussion). We speculate that combining BALF with expensive search methods (e.g., SVD-NAS) could yield excellent results, at the expense of compression time.
@@ 6 Conclusion
We have taken several steps to ensure our results are reproducible. Appendix F provides a detailed description of our experimental setup, and we make our code publicly available at https://github.com/davidgonmar/BALF. The repository includes algorithm implementations, experiment and launch scripts with the exact parameters used, and fixed random seeds for every run. It also provides installation instructions and a complete list of dependencies.
@@ Acknowledgments
I thank Fernando Sancho Caparrini and Miguel Ángel Martínez del Amor for their valuable comments and suggestions, which helped improve this manuscript. I acknowledge the access to computational resources provided by the University of Seville, including the NVIDIA A100 40 GB GPU donated through the NVIDIA Hardware Grant, as well as the RTX 4090 and servers managed by the DeepKnowledge group. I am also grateful for the support of the University of Seville through the Programa de Becas de Iniciación a la Investigación (VIIPPIT-2025-II.1).
Appendix A Appendix Overview
The appendix is organized as follows:
@@ Appendix A Appendix Overview
  • Appendix B expands on the factorization scheme, as well as on the parameter and FLOP counts of layers before and after low-rank factorization. It includes clarification for fully connected, ungrouped convolutional, and grouped convolutional layers and motivates the constraint of using a shared per-group rank for the latter.
@@
  • Appendix C details the construction of uncentered whitening matrices via eigendecomposition, discusses numerical choices and fallbacks, and describes the incremental estimation of second moments and other details.
@@
  • Appendix D discusses the specifics of our rank allocator: it specifies the cost model, and presents the previously mentioned Lagrangian relaxation, together with complexity analysis and experimental results on its behavior.
@@
  • Appendix E states and proves the formal counterpart of Theorem 3 (discussed in the main text), including simplified forms for special cases.
@@
  • Appendix F documents the experimental details of our main experiments, including datasets, models, preprocessing, compression protocol, and other important details.
@@
  • Appendix G includes additional experiments that answer secondary questions, including reports on calibration dataset size ablations, robustness under distribution shift (on CIFAR-10-C), practical speedups, and additional accuracy-complexity curves that did not appear in the main text due to space constraints.
@@
  • Appendix H describes the end-to-end practical implementation of BALF at a high level and provides timing and peak memory measurements of its different parts.
@@
  • Appendix I provides a brief analysis of BALF through the lens of information geometry, following the work of Shumaylov et al. (2025).
@@
  • Appendix J collects complete proofs for results deferred from the main text, namely Theorems 2 and 1 and auxiliary lemmas.
  • Appendix K discloses the role of LLMs in this work.
Appendix B Decomposing the Layers and Complexity After Factorization
@@
In this section, we provide a more thorough discussion of how layers are decomposed. We also provide the parameter counts and FLOP counts before and after decomposition (the latter is used as the cost function in our rank allocator). Given a layer and a rank $P$, our scheme shares the same complexity as basic SVD truncation to rank $P$; the difference is the (expected) quality of the outputs of the factorized layer. We omit biases for simplicity, as these remain constant in the original and factorized layers. $B$ denotes the batch size. Strictly speaking, the FLOP counts reported below should all be doubled (as it stands, we count each multiply-accumulate as one FLOP), but we omit the doubling factor for readability.
B.1 Fully Connected Layers
@@
The total storage needed is the sum of the storage for each matrix, that is, $D_{i}P+PD_{o}$, while the number of FLOPs is $BD_{i}P+BPD_{o}$ (the computation order is algebraically irrelevant, but the order is important for efficiency).
B.2 2D Convolutional Layers
@@
The total number of parameters is $PC_{i}H_{k}W_{k}+C_{o}P$ and the FLOPs are $BH_{o}W_{o}(PC_{i}H_{k}W_{k}+C_{o}P)$. Note that $H_{o}$ and $W_{o}$ subsume the additional convolution hyperparameters (padding, strides, dilation, etc.).
@@
B.3 Grouped 2D Convolutional Layers
@@
with total parameters $C_{o}\frac{C_{i}}{G}H_{k}W_{k}$, and FLOPs $BH_{o}W_{o}C_{o}\frac{C_{i}}{G}H_{k}W_{k}$.
@@
The total parameter storage across all groups is $PC_{i}H_{k}W_{k}+C_{o}P$, and the FLOPs are $B\left(H_{o}W_{o}PC_{i}H_{k}W_{k}+H_{o}W_{o}C_{o}P\right)$.
(The intermediate feature map has $GP$ channels in total—$P$ per group—but the parameter and FLOP counts simplify to the expressions above. As before, applying the (possibly spatial) grouped convolution first, followed by the grouped $1\times 1$, is important for efficiency.)
B.3.1 On the Sharing of Group Ranks
@@
The second reason is that, even if we managed to get an optimal implementation, grouped convolutions are, in essence, $G$ smaller convolutions run in parallel. If we managed to overcome all the implementation issues, it is expected that the largest group would generally be a bottleneck for the others, as the smaller groups would have to wait for them.
Appendix C Obtaining the Uncentered Whitening Matrices
In this section, we provide more details about how the uncentered whitening matrices are obtained. We first begin with a discussion on past work.
The naive option to obtain an uncentered whitening matrix is to use the Cholesky decomposition of the second moments, as explained in (Wang et al., 2025c). An issue is that we cannot compute it in the case of singular positive-semidefinite matrices. In their implementation  [^2] https://github.com/AIoT-MLSys-Lab/SVD-LLM, they address this issue by perturbing the matrix with random noise. Later, in Wang et al. (2025b), the authors note this issue and propose the usage of an SVD-based whitening matrix, which supports data living in a subspace (although their analysis does not seem to cover this). Nevertheless, they show slight end-to-end accuracy gains coming from the SVD-based whitening.
As noted in the main text, we use an eigendecomposition-based scheme, implemented in practice through the $\mathsf{eigh}$ routine in PyTorch (a routine that obtains the eigenvalues and eigenvectors of a Hermitian (symmetric in the case of $\mathbb{R}$) matrix, and is usually faster than calling the corresponding generic eigendecomposition routine). First, we show that the eigendecomposition scheme is a valid approach.
Recall that, for a single group $g$, with ${\bm{X}}=\overline{I}({\bm{\mathsfit{X}}})_{g}\in\mathbb{R}^{N\times I}$ we compute

### v2

@@ 4.4 Putting It All Together: BALF
6:  ${\bm{M}},{\bm{M}}^{+}]\leftarrow\textsc{BatchedUWM}(\frac{1}{N_{l}}\overline{I}_{l}({\bm{\mathsfit{X}}}^{l-1})^{T}\overline{I}_{l}({\bm{\mathsfit{X}}}^{l-1}));\ {\bm{\mathsfit{W}}}\leftarrow\mathcal{M}_{l}.\operatorname{weights}$
@@
13:  ${\bm{U}},{\bm{\Sigma}},{\bm{V}}^{T},{\bm{M}},{\bm{M}}^{+}]\leftarrow F[l]$; ${\bm{W}}_{0}\leftarrow{\bm{M}}{\bm{U}}{\bm{\Sigma}}^{1/2}$; ${\bm{W}}_{1}\leftarrow{\bm{\Sigma}}^{1/2}{\bm{V}}^{T}$
@@
A more detailed statement, its proof, and special cases appear in Appendix D. The tightness of the bound depends on the model’s depth and its weights, and it is typically loose in practical deep networks. Even so, it remains conceptually informative.
@@ 5 Experiments
First, we perform an extensive sweep showcasing the complexity-accuracy curves for FLOP and parameter counts. To highlight which components of BALF are important, each run is performed using different methods (all implemented by us): SVD with the energy criterion, activation-aware SVD with the energy criterion, and BALF (activation-aware decomposition with our rank allocator); see Section 4.3 for an overview.
@@
For ResNet-50, we compare BALF with several pruning methods: DFPC (Narshana et al., 2023), IterTVSPrune (ITVSP) (Murti et al., 2023), and IFM (Chen et al., 2024). For ResNet-20, we compare against the pruning approach FVRCP (He et al., 2020) and the factorization method ALDS (Liebenwein et al., 2021).
@@
Compressing compact models such as MobileNet-V2 is known to be challenging. For this model, we compare against the factorization methods SVD-NAS (Yu & Bouganis, 2023), ALDS (Liebenwein et al., 2021), and LR-S2 (Idelbayev & Carreira-Perpiñán, 2020) (using the no-fine-tuning results for ALDS and LR-S2 reported by Yu & Bouganis (2023)). We also compare with the pruning-restoration method proposed by Lee et al. (2025), using their L2 variant (denoted L2+REST). Note that SVD-NAS requires an expensive search procedure (on the order of hours), and LR-S2 allocates ranks through a training run, making them substantially more expensive than BALF. Table 1 contains the different results. In general, BALF surpasses or is competitive with all the compared methods.
@@
On an RTX 2070 laptop, compressing ImageNet models completes in minutes (under four minutes for all models except ResNeXt-101, which takes less than seven). The most expensive steps can be cached, reducing subsequent compression runs to seconds. We provide a comprehensive benchmark and implementation details in Appendix G.
@@
BALF-F-0.5 50.04-50.04 43.86-43.86 13.84-13.84
BALF-P-0.5 28.18-28.18 50.08-50.08 7.78-7.78
BALF-F-0.7 30.00-30.00 21.69-21.69 3.48-3.48
BALF-P-0.7 13.55-13.55 29.94-29.94 1.96-1.96
BALF-F-0.8 20.00-20.00 9.77-9.77 1.65-1.65
BALF-P-0.8 9.03-9.03 21.92-21.92 1.12-1.12
IFM 5.65-5.65 1.46-1.46
IFM 20.42-20.42 10.45-10.45
ITVSP 4.76-4.76 3.02-3.02
ITVSP 9.98-9.98 10.21-10.21
DFPC 5.66-5.66 5.78-5.78
DFPC 10.92-10.92 13.88-13.88
@@
BALF-F-0.5 50.02-50.02 49.48-49.48 24.89-24.89
BALF-P-0.5 24.75-24.75 49.74-49.74 13.21-13.21
BALF-F-0.7 30.74-30.74 29.10-29.10 6.26-6.26
BALF-P-0.7 12.77-12.77 29.88-29.88 2.50-2.50
BALF-F-0.8 20.03-20.03 17.90-17.90 2.16-2.16
BALF-P-0.8 9.27-9.27 21.70-21.70 1.20-1.20
ALDS 23.71-23.71 1.50-1.50
ALDS 30.05-30.05 3.37-3.37
ALDS 50.00-50.00 25.22-25.22
FVRCP 19.84-19.84 2.5-2.5
FVRCP 29.51-29.51 5.83-5.83
FVRCP 49.35-49.35 24.86-24.86
@@
BALF-F-0.6 40.00-40.00 39.52-39.52 3.88-3.88
BALF-P-0.6 40.36-40.36 39.87-39.87 4.11-4.11
BALF-F-0.7 30.00-30.00 29.64-29.64 1.08-1.08
BALF-P-0.7 30.27-30.27 29.91-29.91 1.09-1.09
BALF-F-0.8 20.03-20.03 19.79-19.79 0.14-0.14
BALF-P-0.8 20.18-20.18 19.94-19.94 0.16-0.16
GTP-15.3 13.07-13.07 $0.00$ $0.00$
GTP-8.8 50.00-50.00 $0.00$ 3.50-3.50
PRACTISE 16.6-16.6 2.5-2.5
DC-ViT 16.6-16.6 0.54-0.54
@@
BALF-P-0.7 2.06-2.06 29.70-29.70 27.65-27.65
BALF-P-0.75 1.11-1.11 24.79-24.79 6.13-6.13
BALF-P-0.8 0.23-0.23 19.80-19.80 0.59-0.59
BALF-P-0.9 0.13-0.13 10.76-10.76 0.25-0.25
BALF-F-0.97 2.51-2.51 1.55-1.55 4.95-4.95
SVD-NAS 15.09-15.09 9.00-9.00 12.54-12.54
ALDS 2.62-2.62 37.61-37.61 16.95-16.95
LR-S2 3.81-3.81 6.24-6.24 17.46-17.46
L2+REST 5.00-5.00 4.42-4.42
L2+REST 10.00-10.00 18.47-18.47
@@
On CIFAR-10-C (Hendrycks & Dietterich, 2019), BALF remains robust to distribution shift: when whitening matrices were estimated on the standard CIFAR-10 training set and evaluation was performed on CIFAR-10-C, the additional accuracy drop was typically less than 2.5 pp. Moreover, we found that a smaller number of calibration samples (about 256 and 1024 for CIFAR-10 and ImageNet, respectively) often yields only marginally worse results than those reported in this section. For instance, see Figure 3 for a sweep on ResNet-50. The remaining results can be found in Appendix F.
@@
Even when surpassing (or being competitive with) other methods, compressing compact models (like MobileNet-V2) remains difficult. End-to-end speedups are notable, but we believe that implementing specialized operators would yield better gains (see Section F.3 for a discussion). We speculate that combining BALF with expensive search methods (e.g., SVD-NAS) could yield excellent results at the expense of compression time.
@@ 6 Conclusion
We have taken several steps to ensure that our results are reproducible. Appendix E provides a detailed description of our experimental setup, and we make our code publicly available at https://github.com/davidgonmar/BALF. The repository includes algorithm implementations, experiment and launch scripts with the exact parameters used, and fixed seeds for every run. It also provides installation instructions and a complete list of dependencies.
@@ Acknowledgments
I thank Fernando Sancho Caparrini and Miguel Ángel Martínez del Amor for their valuable comments and suggestions, which helped improve this manuscript. I acknowledge access to computational resources provided by the University of Seville, including an A100 40GB GPU donated through the NVIDIA Hardware Grant, as well as an RTX 4090, and servers managed by the DeepKnowledge group. I am also grateful for the support of the University of Seville through the Programa de Becas de Iniciación a la Investigación (VIIPPIT-2025-II.1).
Appendix Overview
This appendix is organized as follows:
@@
  • Appendix A expands on the factorization scheme and provides parameter and FLOP counts for layers before and after low-rank factorization. It includes results for fully connected, ungrouped convolutional, and grouped convolutional layers, and motivates the use of a shared per-group rank for the latter.
@@
  • Appendix B details the construction of uncentered whitening matrices via eigendecomposition, discusses numerical choices and fallbacks, and describes the incremental estimation of second moments and other details.
@@
  • Appendix C discusses the specifics of our rank allocator: it specifies the cost model, and presents the previously mentioned Lagrangian relaxation, together with complexity analysis and experimental results on its behavior.
@@
  • Appendix D states and proves the formal counterpart of Theorem 3 (discussed in the main text), including simplified forms for special cases.
@@
  • Appendix E documents the experimental details of our main experiments, including datasets, models, preprocessing, compression protocols, and other important details.
@@
  • Appendix F includes additional experiments that answer secondary questions, including reports on calibration dataset size ablations, robustness under distribution shift (on CIFAR-10-C), practical speedups, and additional accuracy-complexity curves that did not appear in the main text due to space constraints.
@@
  • Appendix G describes the end-to-end practical implementation of BALF at a high level and provides timing and peak memory measurements of its different parts.
@@
  • Appendix H provides a brief analysis of BALF through the lens of information geometry, following the work of Shumaylov et al. (2025).
@@
  • Appendix I collects complete proofs for results deferred from the main text, namely Theorems 2 and 1 and auxiliary lemmas.
  • Appendix J discloses the role of LLMs in this work.
Appendix A Decomposing the Layers and Complexity After Factorization
@@
In this section, we provide a more thorough discussion of how layers are decomposed. We also provide the parameter counts and FLOP counts before and after decomposition (the latter is used as the cost function in our rank allocator). Given a layer and a rank $P$, our scheme has the same complexity as basic SVD truncation to rank $P$; the difference lies in the (expected) quality of the outputs of the factorized layer. We omit biases for simplicity, as these remain constant in the original and factorized layers. $B$ denotes the batch size. Strictly speaking, the FLOP counts reported below should all be doubled (as it stands, we count each multiply-accumulate as one FLOP), but we omit the doubling factor for readability.
A.1 Fully Connected Layers
@@
The total storage needed is the sum of the storage for each matrix, that is, $P(D_{i}+D_{o})$, while the number of FLOPs is $BP(D_{i}+D_{o})$ (the computation order is algebraically irrelevant, but the order is important for efficiency).
A.2 2D Convolutional Layers
@@
The total number of parameters is $P(C_{i}H_{k}W_{k}+C_{o})$, and the FLOPs are $BH_{o}W_{o}P(C_{i}H_{k}W_{k}+C_{o})$. Note that $H_{o}$ and $W_{o}$ subsume the additional convolution hyperparameters (padding, strides, dilation, etc.).
@@
A.3 Grouped 2D Convolutional Layers
@@
with a total of $C_{o}\frac{C_{i}}{G}H_{k}W_{k}$ parameters, and a total of $BH_{o}W_{o}C_{o}\frac{C_{i}}{G}H_{k}W_{k}$ FLOPs.
@@
The total parameter storage across all groups is $P(C_{i}H_{k}W_{k}+C_{o})$, and the FLOPs are $BH_{o}W_{o}P(C_{i}H_{k}W_{k}+C_{o})$.
The intermediate feature map has $PG$ channels in total—$P$ per group—but the parameter and FLOP counts simplify to the expressions above.
A.3.1 On the Sharing of Group Ranks
@@
The second reason is that, even if we managed to achieve an optimal implementation, grouped convolutions are, in essence, $G$ smaller convolutions run in parallel. It is expected that the largest group would generally be a bottleneck for the others, as the smaller groups would have to wait for them.
Appendix B Obtaining the Uncentered Whitening Matrices
In this section, we provide more details about how the uncentered whitening matrices are obtained. We first begin with a discussion of past work.
The naive option to obtain an uncentered whitening matrix is to use the Cholesky decomposition of the second moments, as explained in (Wang et al., 2025c). An issue is that we cannot compute it in the case of singular positive-semidefinite matrices. In their implementation [^2] https://github.com/AIoT-MLSys-Lab/SVD-LLM, they address this issue by perturbing the matrix with random noise. Later, in Wang et al. (2025b), the authors note this issue and propose the use of an SVD-based whitening matrix, which supports data residing in a subspace (although their analysis does not seem to cover this case). Nevertheless, they show slight end-to-end accuracy gains coming from the SVD-based whitening.
As noted in Section 4.2, we use an eigendecomposition-based scheme, implemented in practice through the $\mathsf{eigh}$ routine in PyTorch (a routine that obtains the eigenvalues and eigenvectors of a Hermitian (symmetric in the case of $\mathbb{R}$) matrix, and is usually faster than calling the corresponding generic eigendecomposition routine). First, we show that the eigendecomposition scheme is a valid approach.
Recall that, for a single group $g$, with ${\bm{X}}=\overline{I}({\bm{\mathsfit{X}}})_{g}\in\mathbb{R}^{N\times I}$, we compute